### v2

@@ 3 Results
We first conducted large-scale validation of ECGFounder using an internal test dataset comprising 834,926 ECGs, to assess its consistency with diagnoses confirmed by clinically recognized cardiologists (Figure 2). ECGFounder demonstrated an impressive AUC of at least 0.950 across 82 diagnostic categories and at least 0.900 across 112 categories. Overall, the AUCs were particularly high for categories such as rhythm disorders, conduction abnormalities, cardiac hypertrophy, and myocardial infarction. In contrast, the AUCs were lower for categories involving waveform changes, which may be attributed to the annotation habits of certain physicians who tend to place less emphasis on documenting waveform changes in their diagnoses.
@@
In the internal test set experiments focused on single-lead ECGs, the model demonstrated excellent performance in rhythm-type diagnoses(Figure 2). Notably, it achieved an AUC above 0.95 for common heart rate abnormalities such as normal sinus rhythm, sinus bradycardia, sinus tachycardia, marked sinus bradycardia, sinus arrhythmia, marked sinus arrhythmia, and atrial fibrillation. These diagnoses can be reliably identified from single-lead ECGs. Additionally, the model achieved an AUC above 0.8 for diagnoses including premature ventricular complexes, premature supraventricular complexes, pacemaker, 1st degree AV block, branch block, fascicular block, and chamber enlargement. It is worth noting that for some of these diagnoses, this is the first time such exceptional performance has been observed using single-lead ECGs. In the diagnosis of myocardial infarctions, the model also showed good diagnostic performance for lateral infarct, anterolateral infarct, acute MI, and STEMI. In ECG research, these diagnoses are associated with shifts in the heart’s electrical axis. This suggests that data augmentation methods based on the electrical axis are effective for training single-lead ECG models.
@@
Figure 2: Results of AUROC on 12-lead and single-lead ECG in internal test set.
@@ 4 Discussion
ECGFounder enhances diagnostic performance by learning to identify ECG features associated with cardiovascular diseases, which are typically diagnosed based on specific waveform patterns and rhythm characteristics like the elevated ST segments of myocardial infarction and the irregular fluctuations of atrial fibrillation. These features involve anomalies in cardiac electrical activity, appearing significantly different from normal ECG waveforms. Upon training, ECGFounder can recognize these disease-related waveform patterns and rhythms, accurately diagnosing corresponding cardiovascular conditions. As observed in Table S7 in Supplement and Figure 2, ECGFounder matches or even exceeds the performance of cardiology experts on the internal review test sets. Furthermore, ECGFounder’s sensitivity exceeds that of the average cardiology expert, indicating its ability to more accurately capture subtle signs of cardiovascular diseases that may be overlooked by human experts.
@@ Acknowledgements
Table S6 describes the personal information of 5 cardiologists participating in Cardiologist versus ECGFounder, including working year, gender, age and education level.
@@
S4. Details of Results
@@
To further compare the performance of ECGFounder in clinical diagnosis with that of experts, we compared the performance of the model with that of 5 cardiologists. The results are shown in the figure S2 and tableLABEL:tab:internal.
Table S7: Performance of ECGFounder and cardiologists on the committee test set
ECGFounder Cardiologists
Class Count AUC Sens Spec F1 Sens Spec F1
ANTERIOR INFARCT 2 0.908 1.000 0.906 0.408 0.750 0.999 0.625
ATRIAL FIBRILLATION 59 0.996 1.000 0.972 0.866 0.798 0.981 0.796
ATRIAL FLUTTER 19 0.993 1.000 0.986 0.759 0.568 0.990 0.514
ATRIAL-PACED RHYTHM 3 0.991 1.000 0.988 0.500 0.833 0.995 0.657
INFERIOR INFARCT 1 0.998 1.000 0.998 0.667 1.000 0.997 0.725
LATERAL INFARCT 2 0.998 1.000 0.998 0.667 0.250 0.999 0.250
LEFT BUNDLE BRANCH BLOCK 11 1.000 1.000 0.998 0.909 0.600 0.932 0.371
NORMAL SINUS RHYTHM 364 0.969 0.934 0.926 0.952 0.916 0.862 0.930
RIGHT AXIS DEVIATION 12 0.996 1.000 0.992 0.800 0.594 0.979 0.489
RIGHT BUNDLE BRANCH BLOCK 50 0.984 0.940 0.969 0.847 0.775 0.970 0.762
SINUS BRADYCARDIA 41 0.995 1.000 0.989 0.938 0.717 0.994 0.791
SINUS RHYTHM 191 0.970 0.970 0.926 0.971 0.916 0.862 0.930
SINUS TACHYCARDIA 58 0.996 1.000 0.945 0.683 0.821 0.990 0.833
VENTRICULAR TACHYCARDIA 9 0.903 1.000 0.882 0.635 0.875 0.996 0.678
WITH 1ST DEGREE AV BLOCK 15 0.864 0.667 0.831 0.287 0.633 0.962 0.560
WITH SINUS ARRHYTHMIA 6 0.976 1.000 0.945 0.308 0.542 0.979 0.319
@@
Moreover, as shown in the figure S3 and S4, we evaluated several downstream tasks on single-lead ECGFounder thorough fine-tuning.
@@
S5. Data link
S6. Model weights link
Will be released on https://bdsp.io/
S7. Code link

### v3

@@ 3 Results
We first conducted large-scale validation of ECGFounder using an internal test dataset comprising 834,926 ECGs, to assess its consistency with diagnoses confirmed by clinically recognized cardiologists (Figure S3). ECGFounder demonstrated an impressive AUC of at least 0.950 across 82 diagnostic categories and at least 0.900 across 112 categories. Overall, the AUCs were particularly high for categories such as rhythm disorders, conduction abnormalities, cardiac hypertrophy, and myocardial infarction. In contrast, the AUCs were lower for categories involving waveform changes, which may be attributed to the annotation habits of certain physicians who tend to place less emphasis on documenting waveform changes in their diagnoses.
@@
In the internal test set experiments focused on single-lead ECGs, the model demonstrated excellent performance in rhythm-type diagnoses(Figure S3). Notably, it achieved an AUC above 0.95 for common heart rate abnormalities such as normal sinus rhythm, sinus bradycardia, sinus tachycardia, marked sinus bradycardia, sinus arrhythmia, marked sinus arrhythmia, and atrial fibrillation. These diagnoses can be reliably identified from single-lead ECGs. Additionally, the model achieved an AUC above 0.8 for diagnoses including premature ventricular complexes, premature supraventricular complexes, pacemaker, 1st degree AV block, branch block, fascicular block, and chamber enlargement. It is worth noting that for some of these diagnoses, this is the first time such exceptional performance has been observed using single-lead ECGs. In the diagnosis of myocardial infarctions, the model also showed good diagnostic performance for lateral infarct, anterolateral infarct, acute MI, and STEMI. In ECG research, these diagnoses are associated with shifts in the heart’s electrical axis. This suggests that data augmentation methods based on the electrical axis are effective for training single-lead ECG models.
@@
Table 1: Performance of ECGFounder and cardiologists on the committee test set
ECGFounder Cardiologists
Class Count AUC Sens Spec F1 Sens Spec F1
ANTERIOR INFARCT 2 0.908 1.000 0.906 0.408 0.750 0.999 0.625
ATRIAL FIBRILLATION 59 0.996 1.000 0.972 0.866 0.798 0.981 0.796
ATRIAL FLUTTER 19 0.993 1.000 0.986 0.759 0.568 0.990 0.514
ATRIAL-PACED RHYTHM 3 0.991 1.000 0.988 0.500 0.833 0.995 0.657
INFERIOR INFARCT 1 0.998 1.000 0.998 0.667 1.000 0.997 0.725
LATERAL INFARCT 2 0.998 1.000 0.998 0.667 0.250 0.999 0.250
LEFT BUNDLE BRANCH BLOCK 11 1.000 1.000 0.998 0.909 0.600 0.932 0.371
NORMAL SINUS RHYTHM 364 0.969 0.934 0.926 0.952 0.916 0.862 0.930
RIGHT AXIS DEVIATION 12 0.996 1.000 0.992 0.800 0.594 0.979 0.489
RIGHT BUNDLE BRANCH BLOCK 50 0.984 0.940 0.969 0.847 0.775 0.970 0.762
SINUS BRADYCARDIA 41 0.995 1.000 0.989 0.938 0.717 0.994 0.791
SINUS RHYTHM 191 0.970 0.970 0.926 0.971 0.916 0.862 0.930
SINUS TACHYCARDIA 58 0.996 1.000 0.945 0.683 0.821 0.990 0.833
VENTRICULAR TACHYCARDIA 9 0.903 1.000 0.882 0.635 0.875 0.996 0.678
WITH 1ST DEGREE AV BLOCK 15 0.864 0.667 0.831 0.287 0.633 0.962 0.560
WITH SINUS ARRHYTHMIA 6 0.976 1.000 0.945 0.308 0.542 0.979 0.319
@@ 4 Discussion
ECGFounder enhances diagnostic performance by learning to identify ECG features associated with cardiovascular diseases, which are typically diagnosed based on specific waveform patterns and rhythm characteristics like the elevated ST segments of myocardial infarction and the irregular fluctuations of atrial fibrillation. These features involve anomalies in cardiac electrical activity, appearing significantly different from normal ECG waveforms. Upon training, ECGFounder can recognize these disease-related waveform patterns and rhythms, accurately diagnosing corresponding cardiovascular conditions. As observed in Table S7 in Supplement and Figure S3, ECGFounder matches or even exceeds the performance of cardiology experts on the internal review test sets. Furthermore, ECGFounder’s sensitivity exceeds that of the average cardiology expert, indicating its ability to more accurately capture subtle signs of cardiovascular diseases that may be overlooked by human experts.
@@ Acknowledgements
Table S6 describes the personal information of 4 cardiologists participating in Cardiologist versus ECGFounder, including working year, gender, age and education level.
@@
S3.3. Details of ECG preprocessing
Here, we provide a more comprehensive description of our preprocessing pipeline . The description is as follows:
First, we used linear interpolation to resample each ECG signal to a fixed sampling rate of 500Hz. This step ensures that all ECG signals, regardless of their original sampling rate, have a consistent temporal resolution, facilitating downstream processing and analysis.
Next, we applied a high-pass filter with a cutoff frequency of 1Hz to suppress residual baseline drift. Baseline drift often arises from low-frequency variations such as respiration and electrode-skin impedance changes. By attenuating frequency components below 1Hz, this high-pass filter helps maintain the main morphological features of the ECG waveform without low-frequency noise.
After that, we employed a 2nd-order Butterworth low-pass filter with a cutoff frequency of 30Hz to reduce high-frequency noise. This filter preserves the primary frequency components of the ECG signal—typically within the 0–30Hz range—while attenuating higher frequencies that may be introduced by muscle artifacts or electrical interference.
Furthermore, to remove power-line interference, we used a notch filter at 50/60Hz. This notch filter specifically targets and attenuates the narrow frequency band at which power-line noise occurs, improving signal fidelity without excessively affecting the rest of the ECG spectrum.
Then, from any ECG recording longer than 10 seconds, we extracted consecutive 10-second segments. If a recording was shorter than 10 seconds, we applied zero-padding to the signal to reach the desired 10-second duration. This uniform window length ensures that each segment fed into the model has the same temporal dimension.
Before inputting these processed segments into the model, we performed Z-score normalization using each segment’s mean and standard deviation. Concretely, this involves subtracting the segment mean from each data point and dividing by the segment’s standard deviation, resulting in each segment having a mean of 0 and a standard deviation of 1. This step prevents signals with inherently larger amplitudes from overshadowing those with smaller amplitudes, making model training more stable.
Lastly, for any missing ECG signals, we simply assigned them a value of zero. In other words, these missing channels are represented as zero-filled data, ensuring a consistent input dimension for the model without introducing spurious numerical artifacts.
By following these steps, we effectively mitigate baseline drift, high-frequency noise, and power-line interference, while establishing a uniform sampling rate and window size for all signals, ultimately providing clean and standardized ECG segments for subsequent analysis and model training.
S4. Details of Experiments
S4.1. Details of ablation experiments
We provide a series of ablation experimental results to demonstrate the advantages and innovations of our method.
Here, we use the average AUC of the internal validation set and the average AUC of the CODE-test (external validation set) as our metrics, respectively.
The table are as follow:
Table S7: Effects of loss function.
Loss Internal External
BCE Loss 86.18 (86.12-86.24) 95.32 (95.31-95.33)
Focal Loss 88.67 (88.65-88.69) 96.73 (96.72-96.74)
PU Loss (Ours) 92.04 (92.00-92.07) 98.20 (98.15-98.25)
Table S8: Effects of hyperparameters of loss function.
Num of $gamma$ Internal External
0.5 91.18 (91.14-91.22) 97.22 (97.21-97.23)
1 91.77 (91.75-88.79) 97.37 (97.36-97.38)
1.5 (Ours) 92.04 (92.00-92.07) 98.20 (98.15-98.25)
2 91.96 (91.94-91.98) 98.12 (98.11-98.13)
Table S9: Effects of model architecture.
Model architecture Internal External
CNN-1d 87.45 (87.43-97.47) 93.21 (93.12-93.30)
ResNet-1d 91.85 (91.84-91.86) 97.64 (97.62-97.67)
Transformer-1d 91.53 (91.52-91.55) 97.78 (97.75-97.81)
RegNet-1d (Ours) 92.04 (92.01-92.07) 98.20 (98.18-98.23)
Table S10: Effects of 1-lead data argumentation.
Methods Internal(1-lead) External(1-lead)
Only 1-lead data 83.76 (83.62-83.90) 94.96 (94.83-95.09)
1-lead data argumentation (Ours) 86.92 (86.83-87.01) 96.37 (96.33-96.41)
Table S11: Effects of number of model parameters.
Model architecture Internal External
11.7M 89.62 (89.50-89.74) 96.67 (96.61-96.73)
25.6M 90.23 (90.04-90.41) 97.41 (97.30-97.52)
76.3M (Ours) 92.04 (92.01-92.07) 98.20 (98.18-98.23)
110M 91.87 (91.85-91.88) 98.34 (98.01-98.67)
Table S12: Effects of number of training set.
Model architecture Internal External
10K 86.78 (86.72-86.84) 91.11 (91.04-91.18)
100K 87.93 (87.93-87.93) 92.08 (92.03-92.13)
1M 89.95 (89.73-90.17) 96.21 (96.11-96.31)
3M 91.85 (91.80-91.89) 97.82 (97.72-97.91)
5M 91.93 (91.62-92.25) 97.78 (97.75-97.81)
10M (Ours) 92.04 (92.01-92.07) 98.20 (98.18-98.23)
S4.2. Details of linear probing
In order to explore the advantages of ECGFounder under few-shot fine-tuning and the generalization of features, we conducted linear probing experiments on an ECG self supervised learning benchmark.
For linear probing, we kept the backbone network frozen and only trained the randomly initialized parameters of the linear classifier. To explore the performance of our method under low-resource conditions, we conducted linear probing using 1%, 10%, and 100% of the training data for each task. We set the learning rate to $5\times 10^{-3}$ and trained for 100 epochs. All test results were obtained from the best validation model, rather than testing the model on the test set after each epoch and reporting the highest result. For all downstream tasks, we used the AUC as the evaluation metric. We set the random seed to 00 to ensure the reproducibility of all results.
Table S13: Linear probing results of ECGFounder and other SSL models. The best results are bolded, with gray indicating the second highest.
Method PTBXL-Super PTBXL-Sub PTBXL-Form PTBXL-Rhythm CPSC2018 CSN 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% SimCLR 38 63.41 69.77 73.53 60.84 68.27 73.39 54.98 56.97 62.52 51.41 69.44 77.73 59.78 68.52 76.54 59.02 67.26 73.20 TS-TCC 39 70.73 75.88 78.91 53.54 66.98 77.87 48.04 61.79 71.18 43.34 69.48 78.23 57.07 73.62 78.72 55.26 68.48 76.79 CLOCS 40 68.94 73.36 76.31 57.94 72.55 76.24 51.97 57.96 72.65 47.19 71.88 76.31 59.59 77.78 77.49 54.38 71.93 76.13 ST-MEM 41 61.12 66.87 71.36 54.12 57.86 63.59 55.71 59.99 66.07 51.12 65.44 74.85 56.69 63.32 70.39 59.77 66.87 71.36 HeartLang 42 78.94 85.59 87.52 64.68 79.34 88.91 58.70 63.99 80.23 62.08 76.22 90.34 60.44 66.26 77.87 57.94 68.93 82.49 ECG-FM 11 69.41 82.10 85.79 64.12 79.28 86.51 59.67 66.98 76.02 59.83 74.08 86.05 59.94 69.19 78.81 55.24 66.83 79.05 ECGFounder (Ours) 79.65 87.34 91.20 75.72 81.73 88.03 61.69 68.76 82.52 70.05 88.10 93.61 64.21 79.65 83.18 64.68 75.89 84.35
S4.3. Details of subgroup analysis
In order to explore the fairness of ECGFounder under various demographic conditions, we conducted subgroup analysis on ECGFounder. The table is as follows:
Table S14: Performance in age, sex, and race stratified populations.
Cohort Internal External
Less than 40 years old 91.85 (91.75-91.94) 98.21 (98.18-98.24)
40-60 years old 92.44 (92.39-92.48) 97.94 (97.92-97.96)
60-75 years old 91.58 (91.55-91.62) 97.78 (97.75-97.81)
Greater than 75 years old 92.03 (91.95-92.13) 98.21 (98.20-98.23)
Male 91.98 (91.92-92.14) 98.22 (98.18-98.26)
Female 92.10 (92.06-92.15) 98.16 (98.15-98.17)
White 92.72 (92.66-92.78) /
Asian 91.44 (91.40-92.48) /
African American 91.90 (91.86-91.94) /
American Indian 90.55 (90.46-90.64) /
Other races 92.11 (92.07-92.15) /
S5. Details of Results
S5.1. Details of internal committee validation
@@
To further compare the performance of ECGFounder in clinical diagnosis with that of experts, we compared the performance of the model with that of 5 cardiologists. The results are shown in the figure S2.
Label Sensitivity Sensitivity_CI Specificity Specificity_CI
LATERAL INFARCT 1 (1.0, 1.0) 0.94 (0.92, 0.96)
INFERIOR INFARCT 1 (1.0, 1.0) 1 (0.99, 0.99)
PREMATURE ATRIAL COMPLEXES 0.95 (0.87, 1.0) 0.96 (0.95, 0.97)
SINUS TACHYCARDIA 0.93 (0.85, 1.0) 0.97 (0.96, 0.99)
RIGHT AXIS DEVIATION 1 (1.0, 1.0) 0.97 (0.96, 0.98)
ATRIAL FIBRILLATION 1 (1.0, 1.0) 0.97 (0.95, 0.98)
SINUS BRADYCARDIA 1 (1.0, 1.0) 0.98 (0.97, 0.99)
ATRIAL FLUTTER 0.91 (0.90, 0.94) 0.99 (0.98, 0.99)
ATRIAL-PACED RHYTHM 0.63 (0.32, 0.94) 1 (0.99, 1.0)
RIGHT BUNDLE BRANCH BLOCK 0.92 (0.88, 0.97) 0.97 (0.95, 0.98)
PREMATURE VENTRICULAR COMPLEXES 0.88 (0.79, 0.97) 0.95 (0.93, 0.97)
WITH SINUS ARRHYTHMIA 0.75 (0.67, 0.87) 0.96 (0.94, 0.96)
SINUS RHYTHM 0.98 (0.96, 0.99) 0.54 (0.48, 0.63)
NORMAL SINUS RHYTHM 0.94 (0.91, 0.96) 0.9 (0.89, 0.95)
ANTERIOR INFARCT 0 (0.0, 0.0) 0.93 (0.91, 0.95)
VENTRICULAR TACHYCARDIA 0 (0.0, 0.0) 1 (0.99, 1.0)
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 0.93 (0.85, 1.0) 0.99 (0.98, 1.0)
WITH 1ST DEGREE AV BLOCK 0.66 (0.62, 0.68) 1 (0.99, 1.0)
LEFT BUNDLE BRANCH BLOCK 1 (1.0, 1.0) 0.98 (0.97, 0.99)
VENTRICULAR-PACED RHYTHM 0.94 (0.91, 0.96) 1 (0.99, 1.0)
Label PPV PPV_CI NPV NPV_CI
LATERAL INFARCT 0.03 (0.0, 0.10) 1 (1.0, 1.0)
INFERIOR INFARCT 0.33 (0.0, 0.66) 1 (1.0, 1.0)
PREMATURE ATRIAL COMPLEXES 0.5 (0.39, 0.62) 1 (1.0, 1.0)
SINUS TACHYCARDIA 0.68 (0.51, 0.81) 1 (0.99, 1.0)
RIGHT AXIS DEVIATION 0.35 (0.27, 0.49) 1 (1.0, 1.0)
ATRIAL FIBRILLATION 0.75 (0.67, 0.87) 1 (1.0, 1.0)
SINUS BRADYCARDIA 0.81 (0.67, 0.89) 1 (1.0, 1.0)
ATRIAL FLUTTER 0.63 (0.32, 0.94) 1 (1.0, 1.0)
ATRIAL-PACED RHYTHM 0.68 (0.51, 0.81) 1 (0.99, 1.0)
RIGHT BUNDLE BRANCH BLOCK 0.78 (0.70, 0.88) 0.99 (0.98, 1.0)
PREMATURE VENTRICULAR COMPLEXES 0.63 (0.54, 0.74) 0.99 (0.98, 1.0)
WITH SINUS ARRHYTHMIA 0.15 (0.05, 0.25) 1 (0.99, 1.0)
SINUS RHYTHM 0.85 (0.81, 0.88) 0.91 (0.86, 0.98)
NORMAL SINUS RHYTHM 0.96 (0.95, 0.97) 0.84 (0.80, 0.92)
ANTERIOR INFARCT 0 (0.0, 0.0) 1 (1.0, 1.0)
VENTRICULAR TACHYCARDIA 0 (0.0, 0.0) 1 (0.99, 1.0)
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 0.13 (0.0, 0.34) 0.99 (0.99, 1.0)
WITH 1ST DEGREE AV BLOCK 0.5 (0.0, 1.0) 0.97 (0.96, 0.99)
LEFT BUNDLE BRANCH BLOCK 0.33 (0.10, 0.51) 1 (1.0, 1.0)
VENTRICULAR-PACED RHYTHM 0.71 (0.71, 1.0) 0.97 (0.96, 0.98)
Label AUROC AUROC_CI AUPRC AUPRC_CI
LATERAL INFARCT 1 (1.0, 1.0) 0.62 (0.57, 0.67)
INFERIOR INFARCT 1 (1.0, 1.0) 0.5 (0.39, 0.61)
PREMATURE ATRIAL COMPLEXES 1 (0.99, 1.0) 0.94 (0.86, 0.98)
SINUS TACHYCARDIA 1 (0.99, 1.0) 0.96 (0.90, 0.99)
RIGHT AXIS DEVIATION 1 (0.99, 1.0) 0.71 (0.29, 0.95)
ATRIAL FIBRILLATION 1 (0.99, 1.0) 0.95 (0.87, 0.98)
SINUS BRADYCARDIA 1 (0.99, 1.0) 0.9 (0.83, 0.99)
ATRIAL FLUTTER 0.99 (0.99, 1.0) 0.62 (0.34, 0.94)
ATRIAL-PACED RHYTHM 0.99 (0.99, 1.0) 0.35 (0.27, 0.49)
RIGHT BUNDLE BRANCH BLOCK 0.98 (0.97, 0.99) 0.81 (0.65, 0.89)
PREMATURE VENTRICULAR COMPLEXES 0.98 (0.97, 0.99) 0.81 (0.71, 0.91)
WITH SINUS ARRHYTHMIA 0.98 (0.96, 1.0) 0.37 (0.11, 0.49)
SINUS RHYTHM 0.97 (0.96, 0.99) 0.98 (0.95, 1.0)
NORMAL SINUS RHYTHM 0.97 (0.94, 0.98) 0.98 (0.97, 0.99)
ANTERIOR INFARCT 0.91 (0.86, 0.98) 0.66 (0.62, 0.68)
VENTRICULAR TACHYCARDIA 0.9 (0.89, 0.95) 0.54 (0.48, 0.63)
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 0.88 (0.79, 0.97) 0.33 (0.10, 0.51)
WITH 1ST DEGREE AV BLOCK 0.86 (0.80, 0.92) 0.22 (0.13, 0.41)
LEFT BUNDLE BRANCH BLOCK 1 (1.0, 1.0) 1 (1.0, 1.0)
VENTRICULAR-PACED RHYTHM 0.99 (0.98, 1.0) 0.74 (0.58, 0.89)
Table S15: The evaluation results of each cardiologist.
Cardiologist 1 Cardiologist 2
Class Sens Spec F1 Sens Spec F1
ANTERIOR INFARCT 1.000 1.000 1.000 1.000 0.996 0.500
ATRIAL FIBRILLATION 0.810 0.969 0.756 0.929 0.991 0.918
ATRIAL FLUTTER 0.818 0.982 0.521 0.909 0.992 0.544
ATRIAL-PACED RHYTHM 1.000 0.996 0.750 1.000 0.990 0.545
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 0.750 0.938 0.248 0.250 0.998 0.333
INFERIOR INFARCT 1.000 0.996 0.500 1.000 0.994 0.400
LATERAL INFARCT 0.000 1.000 0.000 0.000 0.998 0.000
LEFT BUNDLE BRANCH BLOCK 0.454 0.949 0.125 0.727 1.000 0.189
NORMAL SINUS RHYTHM 0.929 0.912 0.947 0.959 0.897 0.960
PREMATURE ATRIAL COMPLEXES 0.368 0.988 0.438 0.842 0.992 0.821
PREMATURE VENTRICULAR COMPLEXES 0.830 0.993 0.868 0.930 1.000 0.961
RIGHT AXIS DEVIATION 0.380 0.996 0.482 0.250 1.000 0.498
RIGHT BUNDLE BRANCH BLOCK 0.500 0.942 0.495 0.920 0.989 0.911
SINUS BRADYCARDIA 0.737 0.998 0.736 0.816 0.996 0.873
SINUS RHYTHM 0.929 0.912 0.947 0.959 0.897 0.960
SINUS TACHYCARDIA 0.929 0.998 0.945 0.714 0.998 0.816
VENTRICULAR TACHYCARDIA 1.000 0.998 0.667 0.500 1.000 0.566
VENTRICULAR-PACED RHYTHM 0.950 0.994 0.705 1.000 0.994 0.730
WITH 1ST DEGREE AV BLOCK 0.600 0.994 0.667 0.867 0.994 0.639
WITH SINUS ARRHYTHMIA 0.000 0.998 0.167 0.833 0.990 0.517
Table S16: The evaluation results of each cardiologist.
Cardiologist 3 Cardiologist 4
Class Sens Spec F1 Sens Spec F1
ANTERIOR INFARCT 1.000 1.000 1.000 0.000 1.000 0.000
ATRIAL FIBRILLATION 0.738 0.991 0.805 0.714 0.974 0.714
ATRIAL FLUTTER 0.091 0.998 0.409 0.455 0.992 0.493
ATRIAL-PACED RHYTHM 0.667 1.000 0.657 0.667 0.998 0.667
INCOMPLETE RIGHT BUNDLE BRANCH BLOCK 1.000 0.996 0.286 0.250 0.996 0.286
INFERIOR INFARCT 1.000 1.000 1.000 1.000 1.000 1.000
LATERAL INFARCT 0.000 1.000 0.000 1.000 1.000 1.000
LEFT BUNDLE BRANCH BLOCK 0.727 0.998 0.800 0.454 0.994 0.400
NORMAL SINUS RHYTHM 0.970 0.934 0.967 0.808 0.963 0.887
PREMATURE ATRIAL COMPLEXES 0.474 0.990 0.684 0.632 0.988 0.649
PREMATURE VENTRICULAR COMPLEXES 0.830 0.991 0.872 0.830 1.000 0.904
RIGHT AXIS DEVIATION 0.880 1.000 0.669 0.880 0.996 0.534
RIGHT BUNDLE BRANCH BLOCK 0.880 0.996 0.805 0.800 1.000 0.889
SINUS BRADYCARDIA 0.447 0.998 0.632 0.868 0.985 0.886
SINUS RHYTHM 0.970 0.934 0.967 0.808 0.963 0.887
SINUS TACHYCARDIA 0.750 0.994 0.852 0.893 0.970 0.746
VENTRICULAR TACHYCARDIA 1.000 1.000 0.800 1.000 0.998 0.667
VENTRICULAR-PACED RHYTHM 0.200 0.998 0.885 0.750 1.000 0.657
WITH 1ST DEGREE AV BLOCK 0.467 0.994 0.560 0.600 0.984 0.563
WITH SINUS ARRHYTHMIA 0.500 0.990 0.311 0.833 0.941 0.250
@@
newpage
S5.2. Details of internal validation
Figure S3: Results of AUROC on 12-lead and single-lead ECG in internal test set.
S5.3. Details of external validation
@@
S5.4. Details of fine-tuning
Moreover, as shown in the figure S4 and S5, we evaluated several downstream tasks on single-lead ECGFounder thorough fine-tuning.
@@
Also, as shown in the fig S6, we compare the performance of linear probing versus full fine-tuning, particularly focusing on the downstream classification tasks of CHD and LVEF. The detailed results clearly demonstrate that full fine-tuning consistently achieves better performance compared to linear probing, particularly for the ECGFounder model. Furthermore, the ECGFounder model shows superior performance compared to SIMCLR and random initialization under both fine-tuning methods, underscoring the strong representational capacity of ECGFounder.
Figure S6: Comparison of different fine-tuning methods, taking the classification of CHD and LVEF as examples.
Table S18: The PPV and prevalence of the fine-tuning task.
Task PPV(%) Prevalence(%)
Age 77.1 53.9
CHD 65 41.4
CKD 24.2 13.2
Lab 94.1 82.4
LVEF 94.1 76.9
Sex 83.7 51.7
Figure S7: 18 images arranged in a 3x6 grid.
S6. Data link
S7. Model weights link
S8. Code link